OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Show how to solve this using any method.

OpenStudy (wasiqss):

x+y=17= x-y=26 solve these equatios to get

OpenStudy (wasiqss):

i mean x+y=170

OpenStudy (anonymous):

solve the equations to get....

OpenStudy (anonymous):

?

OpenStudy (wasiqss):

to get value of two integers

OpenStudy (anonymous):

thank you:)